OpenStudy (anonymous):

The sum of two numbers is 42. The larger number is 8 more than the smaller number. what are the numbers?

OpenStudy (anonymous):

actually needa learn this 😭

OpenStudy (compassionate):

x + y = 42, right? The larger number is 8 more than the smaller number (42 - larger number) = smaller number Now, the larger number is 8 more than the smaller number (42 - larger number) = smaller number + 8 (42 - x) = y + 8

OpenStudy (anonymous):

and solve that?

OpenStudy (compassionate):

I'm not your work horse. You solve it or ask for help on how to solve it.

OpenStudy (anonymous):

im not telling you to do it? im asking i solve that? and thats the answer?

OpenStudy (compassionate):

I must go. @Hero can help you

OpenStudy (compassionate):

You need to isolate x and y then plug-in their repsective terms

OpenStudy (anonymous):

meaning? 😁😁

hero (hero):

Yes, you can solve it @mattrobakk

OpenStudy (anonymous):

-x=y-34

hero (hero):

but define the variables properly Let x be the smaller and y be the larger number. Then x+y=42 y= x+8

OpenStudy (jhannybean):

Solve it as a system of 2 equations, perhaps

OpenStudy (anonymous):

get it to y=x+8 and then?

OpenStudy (anonymous):

plug it in?

OpenStudy (anonymous):

17 and 25?

hero (hero):

Correct

OpenStudy (anonymous):

thanks a bunch!! have one more hard question ...can you guys stay and help?